Show that each of the following equations are dimensionally correct or in correct: x=distance, v=velocity, a=acceleration, t=time

A) x= v^2/4a

b) v=1/3πR^3h

where V=volume of conical shape, π=contant, R=radius, and h=height
Answer #1

A) LHS dimension of x = [L]

RHS v^2 / 4a = [LT^-1]^2 / [LT-2]

= [L^2-1 T^-2+2]

= [L]

Hence LHS = RHS , Correct

b) LHS = [L^3]

RHS = R^3H = [L^3][L] = [L^4]

LHS is not equal to RHS Hence Incorrect